40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 324 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  588   Sat Jun 28 14:56:44 2008 JohnUpdateComputersini files
In short, I was editing the ini files yesterday evening, I didn't e-log it and after some investigation this afternoon it apears
that I am to blame for all the computer problems which followed.


I wanted to edit C0EDCU.ini and C1IOOF.ini to change C1: PSL-FSS_FAST to a fast channel as C1: PSL-FSS_FAST_F
was dead.

I opened these files and made backups. It appears this is where it all went awry. My backup for C1IOOF
is called C1IOO.ini.090627 i.e. missing the F.

Later c1susvme2 and c1iovme crashed. After failing to bring c1iovme back I wondered if my edits had
caused the problems so I restored the back up files. It appears that here I wrote over C1IOO with my backup of
C1IOOF (presumably because I had made a typo in the name).

To remedy the situation we could restore C1IOO from e.g. chans/archive/C1IOO_080618_160028.ini

No excuses for not e-logging this activity.
  587   Sat Jun 28 03:10:25 2008 robUpdateComputersc1iovme

Quote:
C1susvme2 and C1iovme crashed which sent the optics swinging and tripped the watchdogs.

Koji and I were able to restore c1susvme2 without any trouble.

We have been unable to revive c1iovme. We have tried telneting in and running startup.cmd,
the process runs for a while then hangs with "DAQ init failed -- exiting".

Resetting the board doesn't help. I didn't try keying the whole crate.

All optics are back to normal with damping restored.


I tried keying the crate, then keying the DAQ controller & AWG, then powering down & restarting the framebuilder.
On coming up, the framebuild doesn't start a daqd process, and I can't get one to start by hand (it just prints "652", and then stops).
No error messages and daqd doesn't appear in the prstat.

I then tried keying the DAQ controller again (after the fb0 reboot), which blew the watchdogs on all the suspensions. So then I went around and keyed all the crates.

Now, the suspension controllers are back online. Still no c1iovme, and now the framebuilder/DAQ/AWG are also hosed. We can try keying all the crates again, in the order that Yoichi did last week.

After some more poking around, I found the daqd log file. It's now complaining about

Jun 28 03:00:39 fb daqd[546]: [ID 355684 user.info] Fatal error: channel `C1: PSL-FSS_MIXERM_F' is duplicated 126

This is the second error message like this. It first complained about C1: PSL-FSS_FAST_F, so I commented that out of C1IOOF.ini and rebooted the framebuilder (note this is an actual reboot of the full solaris machine). Eventually I discovered that C1IOOF.ini and C1IOO.ini are essentially identical. They presumably will keep getting these duplicate channel errors until one of them is completely removed.

C1IOO.ini has a modification time of seven PM on Friday night. Who did this and didn't elog it? I've now modified C1IOOF.ini, and I don't remember when it was last modified.
  586   Fri Jun 27 19:59:44 2008 JohnUpdateComputersc1iovme
C1susvme2 and C1iovme crashed which sent the optics swinging and tripped the watchdogs.

Koji and I were able to restore c1susvme2 without any trouble.

We have been unable to revive c1iovme. We have tried telneting in and running startup.cmd,
the process runs for a while then hangs with "DAQ init failed -- exiting".

Resetting the board doesn't help. I didn't try keying the whole crate.

All optics are back to normal with damping restored.
  585   Fri Jun 27 18:21:01 2008 JenneUpdateElectronicsResponse of the LO input on the MC demod board
The alarm handler has been flipping out saying that the LO input of the MC's demod board is too low, so at Rana's suggestion, Eric and I measured the response of the LO input. We used an SR345 function generator at 29.485MHz and several different amplitudes to make a table. The demod board should see an input from the LO between 0-2dBm. When I measured what was going into the LO input from the 29.5MHz delay line phase shifter, the LO input was seeing 4dBm. I'm going to put a 3dB attenuator between the phase shifter and the demod board.

Also, now that we have this table of response values, I'm going to change the settings of the alarm handler to something more reasonable.
Amplitude of 29.485MHz input sine wave [dBm]    |        Value of channel C1:IOO-MC_DEMOD_LO
--------------------------------------------    |        -----------------------------------
-10                                             |        -0.000449867
-8                                              |        -0.000449867
-6                                              |        -0.000449867
-4                                              |        0.000384331
-2                                              |        0.00526733
0                                               |        0.0199163
2                                               |        0.0492143
4                                               |        0.0931613
6                                               |        0.161523
8                                               |        0.229885
10                                              |        0.293364
  584   Fri Jun 27 18:03:46 2008 JenneUpdateElectronicsAnother bad cable in the MC servo
Eric was helping me to measure the response of the LO input on the MC's Demod board, and when we disconnected the end of the cable between the demod board and the delay line phase shifter for the 29.5MHz oscillator, we noticed that the phase shifter's end of the cable was loose, like the connector wasn't fully connected. When we checked it by wiggling the connector, the SMA end fell off. I made a new SMA end for the cable, and reinstalled the cable. The MC locked as soon as I plugged the cable in, so everything seems good again. I tried to not change the cable length when I remade the connector, but the cable is shorter than it was by a small amount, due to the way the end fell off.
  583   Fri Jun 27 15:20:52 2008 robDAQLSC.ini file change

I removed C1:LSC-XARM_CTRL from the frames and added C1:LSC-CARM_ERR
  582   Fri Jun 27 14:36:39 2008 JohnSummaryLocking 
Rob, Yoichi, John

Some progress last night:

Switched back to SPOB_DC for CARM.

Shaped the MC LSC loop to reduce excess noise in the 20-30Hz band. Likely the most significant change.
Could this be due to fan noise from the laptop on the SPOB table?

Brought in the AO path earlier (at low gain).

Reduced offset to 6 and increased MC gain before handing off to SPOB. Ramped up AO and MC gain then switched off the
moving zero.

Looks like PD11 is the most promising candidate for RF CARM although the demod phase needs attention.
Attachment 1: gettingcloser.png
gettingcloser.png
  581   Fri Jun 27 09:20:15 2008 steveFrogsPEMdust particle count is up & alarm handler is on
This 3 years plot show the trend of seasons.
When outside air quality goes bad ( 0.5 micron > 1 million ) the lab follows.
I will demonstrate this effect with a 4th of July fire works calibration.
Let's do not forget the construction activity next door either.

The alarm handler is busy:
It's sound level were reset to a modest level yesterday.
It would be nice to change the alarm sound so it can play Wagner:Der Ring des Nibelungen:
Das Rheingold, Die Walkure & Siegfried and Gotterdammerung
or something more appropriate than the present frog call

1, half micron count is climbing ( it's close to 16k now )
2, MZ refl signal is too high
3, MC lenght servo LO is too
Attachment 1: dust3y.jpg
dust3y.jpg
  580   Thu Jun 26 22:08:33 2008 JenneUpdateElectronics3.7MHz bandstop filter in MC Servo
The 3.7MHz elliptical bandstop filter that I made during my SURF summer is now installed in the MC servo loop to reduce the noise at 3.7MHz.

I have taken transfer functions with and without the filter between TP1A and TP2A, with EXCA at -20dBm, using the HP4195A Network Analyzer. I have also taken power spectra of TP1A with and without the filter, and time domain data with the filter of OUT2 on the MC Servo Board and Qmon on the Demod board just before the MC servo board. The filter is between Qmon and OUT2 in the loop.

The UGF and phase margin don't change noticeably with and without the filter, and the MC still locks nicely (after the minor fiasco this afternoon), so I think it's okay. The UGF is around 57kHz, with about 38 degrees phase margin.

1 July 2008: I redid the plots. Same info, but the traces with and without the filters are now on the same graph for easier readability.
Attachment 1: MCLoopGainBoth.png
MCLoopGainBoth.png
Attachment 2: TP1ASpectrumBoth.png
TP1ASpectrumBoth.png
Attachment 3: QmonWithFilt.png
QmonWithFilt.png
Attachment 4: MCOut2WithFilt.png
MCOut2WithFilt.png
  579   Thu Jun 26 21:07:11 2008 ranaConfigurationASSdust & MC1
I realized today, that yesterday while we were trying to debug the adaptive noise canceler, I turned
off the analog dewhitening on MC1. I did this by changing settings on the Xycom screen but I
forgot to elog this -- this may have caused problems with locking via increased frequency noise.
I have now returned it to its nominal, dewhitening on, configuration.

I also used mDV to look at the last year of dust trend. I have plotted here the cumulative
histogram in percentile units of the 0.5 micron dust level. The x-axis is in units of particles per cu. ft.
and the y-axis is percentage. For example, the plot tells us that over the last year, the counts were
below 6000, 90% of the time. I have set the yellow and red alarm levels to alarm at the 95-th and 99-th
percentile levels, respectively.
Attachment 1: Screenshot-2.png
Screenshot-2.png
  578   Thu Jun 26 20:59:22 2008 ranaUpdateComputer Scripts / ProgramsAlarm Handler
Please do not turn off the Alarm Handler on op540m.
  577   Thu Jun 26 18:29:44 2008 JenneUpdate MC Back online
Jenne, Rob, Yoichi

I was playing with the Mode Cleaner earlier today, working on measuring the effect of the new filter (see next post for loop gain measurements etc.), and bumped something which made it so the Mode Cleaner would not lock.

After much poking around by Rob and Yoichi, we found that the TNC-to-2 pin LEMO from Out1 of the MC Servo Board to Ch1 of the Pentek Generic board to the right of the MC Servo Board was bad. If we touched the cable, the MC would lock, but as soon as we let go, the MC would fall out of lock. The LEMO end of the cable was not heat shrink wrapped, so the 2 wires could have been touching. I replaced the cable, and the MC locked immediately after plugging it in, so I think that has fixed the problem.
  576   Thu Jun 26 18:27:04 2008 ranaUpdateComputer Scripts / ProgramsNo Dataviewer No Cry
Dataviewer was hanging. It would open the Grace window but then not plot anything. Since DTT was working
fine we diagnosed this as a DV only problem. It turned out that there was a boatload of messages in the
dataviewer queue. You can check for extra messaages using the command 'ipcs -q' and then use 'rmq' to remove them.

Dataviewer is working now.

Here's the screen dump from op440m:
op440m:~>ipcs -q
IPC status from <running system> as of Thu Jun 26 18:19:32 PDT 2008
T         ID      KEY        MODE        OWNER    GROUP
Message Queues:
q        400   0x1a51     -Rrw-rw-rw- controls    staff
q          1   0x501      -Rrw-rw-rw- controls    staff
q          2   0x512      -Rrw-rw-rw- controls    staff
q        103   0x553      -Rrw-rw-rw- controls    staff
q          4   0x574      -Rrw-rw-rw- controls    staff
q          5   0x18be     --rw-rw-rw- controls    staff
q         56   0x1d9f     -Rrw-rw-rw- controls    staff
q          7   0x1e56     -Rrw-rw-rw- controls    staff
q          8   0x1efd     -Rrw-rw-rw- controls    staff
q        109   0x2044     -Rrw-rw-rw- controls    staff
q         10   0x207a     -Rrw-rw-rw- controls    staff
q         11   0x24c4     -Rrw-rw-rw- controls    staff
q         12   0x24d7     -Rrw-rw-rw- controls    staff
q         13   0x251b     -Rrw-rw-rw- controls    staff
q         14   0x2fe8     -Rrw-rw-rw- controls    staff
q        115   0x3a65     -Rrw-rw-rw- controls    staff
q        116   0x967      -Rrw-rw-rw- controls    staff
q         17   0x98e      -Rrw-rw-rw- controls    staff
q        118   0x456c     -Rrw-rw-rw- controls    staff
q        669   0x194      -Rrw-rw-rw- controls    staff
q        620   0xe70      -Rrw-rw-rw- controls    staff
q         71   0x15d0     -Rrw-rw-rw- controls    staff
q        222   0x3586     -Rrw-rw-rw- controls    staff
q        173   0x5e37     -Rrw-rw-rw- controls    staff
q        624   0x5e4d     -Rrw-rw-rw- controls    staff
q        475   0x10d7     -Rrw-rw-rw- controls    staff
q        176   0x5ee3     -Rrw-rw-rw- controls    staff
q        277   0x4e22     -Rrw-rw-rw- controls    staff
q        428   0x25f3     -Rrw-rw-rw- controls    staff
q         29   0x3354     -Rrw-rw-rw- controls    staff
q        180   0x368f     -Rrw-rw-rw- controls    staff
q        131   0xb0f      -Rrw-rw-rw- controls    staff
q         82   0x47a5     --rw-rw-rw- controls    staff
q         83   0x4b83     -Rrw-rw-rw- controls    staff
q         34   0x4f6d     -Rrw-rw-rw- controls    staff
q         85   0x4539     -Rrw-rw-rw- controls    staff
q         86   0x67a1     --rw-rw-rw- controls    staff
q         37   0x4b7      -Rrw-rw-rw- controls    staff
q         38   0xd37      -Rrw-rw-rw- controls    staff
q         39   0x156c     -Rrw-rw-rw- controls    staff
q         40   0x2b62     -Rrw-rw-rw- controls    staff
op440m:~>rmq
---------------------------------------------------------------
Deleting message queues which are owned by user: controls ....
---------------------------------------------------------------
      Deleted message queue: 400
      Deleted message queue: 1
      Deleted message queue: 2
      Deleted message queue: 103
      Deleted message queue: 4
      Deleted message queue: 5
      Deleted message queue: 56
      Deleted message queue: 7
      Deleted message queue: 8
      Deleted message queue: 109
      Deleted message queue: 10
      Deleted message queue: 11
      Deleted message queue: 12
      Deleted message queue: 13
      Deleted message queue: 14
      Deleted message queue: 115
      Deleted message queue: 116
      Deleted message queue: 17
      Deleted message queue: 118
      Deleted message queue: 669
      Deleted message queue: 620
      Deleted message queue: 71
      Deleted message queue: 222
      Deleted message queue: 173
      Deleted message queue: 624
      Deleted message queue: 475
      Deleted message queue: 176
      Deleted message queue: 277
      Deleted message queue: 428
      Deleted message queue: 29
      Deleted message queue: 180
      Deleted message queue: 131
      Deleted message queue: 82
      Deleted message queue: 83
      Deleted message queue: 34
      Deleted message queue: 85
      Deleted message queue: 86
      Deleted message queue: 37
      Deleted message queue: 38
      Deleted message queue: 39
      Deleted message queue: 40
---------------------------------------------------------------
op440m:~>
  575   Thu Jun 26 18:24:28 2008 YoichiUpdatePSLFSS input ofset slider problem - fixed
I checked the FSS servo interface board and found that a LT1125CSW used to differentialize offset channel was broken (no virtual short).
So I replaced it. Now the slider is working.
The op-amp was hitting the rail. So it seems like we had been applying the maximum offset to the FSS input all the time.
The reason why the FSS loop still worked with the large offset is that the applied offset (~14V) is attenuated by a factor of 500 at the summing point.
  574   Thu Jun 26 14:06:00 2008 MashaUpdateGeneral500mW INNOLIGHT NPRO info
Below is the placement of 500mW INNOLIGHT NPRO mephisto laser. It is set up on the Symmetric Port table.
  573   Thu Jun 26 12:30:40 2008 JohnSummaryLockingCARM on REFL_DC
Idea:

Try REFL_DC as the error signal for CARM rather than PO_DC.

Reasoning:

The PO signal is dominated by sideband light when the arms are detuned so that any misalignment in the recycling cavity introduces spurious signals. Also, the transfer function from coupled cavity excitation to REFL signal is not so steep and hence REFL should give a little more phase. Finally, the slope of the REFL signal should make it easier to hand over to RF CARM.

Conclusion:


The REFL signal showed no clear improvement over PO signals. We've gone back to PO.


During the night we also discovered that the LO for the MC loop is low.
  572   Thu Jun 26 10:56:15 2008 Max JonesUpdatePEMRemoved Magnetometer
I removed the Bartington Magnetometer on the x arm to one of the outside benches. I'll be trying to determine if and how it works today. It makes a horrible high pitched sound which is due to the fact that the battery is probably 16 yrs old. It still works with ac power though and I want to see if it is still operating correctly before I ask to buy a new battery. Sorry for the bother.
  571   Thu Jun 26 01:10:18 2008 ranaUpdatePEMAlarm Handler indicates that dust level is high
In its first useful act, the Alarm Handler started beeping indicating that the dust particle
counts for particles of diameter less than 0.5 micron had exceeded 5000 /cu. ft.
Here's the
80 day trend of particles, temperature and humidity:
Attachment 1: Untitled.png
Untitled.png
  570   Thu Jun 26 01:08:22 2008 ranaConfigurationGeneralAlarm Handler Revived
I have revived the Alarm Handler by turning it on on op540m and adjusting the levels of
several of the alarming channels to not alarm (like laser power). The alarm levels are now
set to something reasonable and people should start actually paying attention to them.

I also removed the EO Shutter and Stacis alarm stuff since we don't use them.

To really get in and edit it, you have to close the Alarm Handler and edit the file
in /cvs/cds/caltech/alh/. It allows you to add/subtract useful channels and put in
guidance information.

If the alarm handler beeps about something, don't just close it or silence it, Steve. Just
fix it somehow (either set the threshold better or find the real cause).
Attachment 1: b.gif
b.gif
  569   Wed Jun 25 18:03:21 2008 YoichiConfigurationPSLFSS Input Offset slider problem
While working on the PMC scanning, I noticed that the FSS input offset slider is doing nothing.
I traced the signal flow and checked the cables/boards.
The slider changes the output voltage from a VMIVME4116 DAC in the PSL rack. This output voltage is confirmed to be correct at the FLKM64 connector. The signal is connected to the FSS servo interface box (D040423) trough a ribbon cable. However, the output from the interface box is always -27V regardless of the slider position.
Therefore, either the interface box (D040423) or the ribbon cable has a problem.
I will debug the interface box using an extension card when no one is working on the interferometer.
  568   Wed Jun 25 13:56:56 2008 JohnSummaryLockingTuesday night locking
Rob, John

Worked to try and reduce the CARM offset using the dc arm transmissions before changing to SPOB DC. This proved somewhat unsuccessful, the offset couldn't be reduced to more than five (arms storing 5x more power than single arm cavity lock).
  567   Wed Jun 25 13:38:22 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Placement of the 700mW NPRO on the AP table
This morning I have put the 700mW NPRO on the AP table for the abs length measurement.

The RF amplifier was moved (the cables were not changed). I cleaned up some cable arrangements. I was keen not to disturb any of the other optical path. Even so, please let me know if any suspicious behaviour is found on the AP table.
Attachment 1: NPRO700mW_placement.jpg
NPRO700mW_placement.jpg
  566   Wed Jun 25 12:25:28 2008 EricSummaryCameras2D Gaussian Fitting Code
I initially wrote a script in MATLAB that takes pictures of the laser beam's profile and fits them to a two dimensional gaussian in order to determine the position and width of the beam. This code is now (mostly) ported to C so that it can be imbedded in the camera software package that Joe is writing. The fitting works fairly well for pictures with the beam directly incident on the camera, and less well for pictures of scatter off the end mirrors of the arms, since scatter from defects in the mirror have intensities much greater than the intensity of the beam's gaussian profile.

The next steps are to finish up porting the fitting code to C, and then modify it so it can better handle the images off the end mirror. Some thoughts on how to do this are to use a fourier transform and a low pass filter, or to simply use a center-of-mass calculation (with the defect peaks reduced in intensity), since position is more important than beam width in this calculation. The eventual goal is to include the edge of the optic in the picture and use the fit of the beam position in comparison to the optic's position to find the beam's location on the mirror.
  565   Wed Jun 25 11:36:14 2008 Max JonesUpdateComputer Scripts / ProgramsFirst Week Update
For the first week I have been modifying the noise budget script in caltech/NB to run with 40 m parameters and data. As per Rana's instructions, I have tried to run the script with only seismic and Darm sources. This involves identifying and changing channel names and altering paramter files (such as NB/ReferenceData/C1IFOparams.m). To supply the parameter files, I have copied the H1 files with (as yet only) slight modification. The channel name changes have been made to mirror the sites for the most part. Two figures are attached which show the current noise budget. The Day plot was taken 6/23/08 at ~10:30 am. The night plot was taken 6/22/08 at ~11:00 pm . Note that the SRD curve is for the sites and not for the 40 m (I hope to change that soon). Also in one of the plots the DARM noise signal is visible. Obviously this needs work. A list of current concerns is

1) I am using a seismic transfer function made by previous SURF student Ryan Kinney to operate with channels of the form C1:PEM_ACC-ETMY_Y (should I be using C1:DMF-IX_ACCY?) and the channels I am currently using are the acceleraometers for the mode cleaner with names of the form C1:PEM_ACC-MC1_X. Rana said that he thinks these may be the same but I need to be sure.

2) We don't have a DARM_CTRL channel but the code requires it, currently I am using DARM_ERR as a substitute which is probably partly responsible for the obvious error in DARM noise.

Any suggestions are appreciated. Max.
Attachment 1: C1_NoiseBudgetPlot_Day.eps
C1_NoiseBudgetPlot_Day.eps
Attachment 2: C1_NoiseBudgetPlot_Night.eps
C1_NoiseBudgetPlot_Night.eps
  564   Wed Jun 25 11:01:45 2008 MashaUpdateAuxiliary lockingfiber stabilization
For the first week, I have been learning about fiber noise cancellation, auxiliary locking techniques, and other relevant helpful topics in more detail. I am now working on a setup (more detail in previous entry) to measure phase noise introduced by 25m(?) fiber, and then will proceed to try to cancel the noise.
  563   Wed Jun 25 09:46:45 2008 SharonUpdate Adaptive Filters
I have been learning about different methods for applying adaptive filters to improve the Mode Cleaner lock in specific, and other LIGO systems in general.
Finding the exact number of coeffs we would like to apply for our FIR adaptive filter is very important to the efficiency of the filter. Getting this number higher might improve the accuracy of the filter, but costs time we do not have. Another important number to find is the step size. The step size is the variable that controls how far back we want to look into our data for finding the new coeffs. To understand more about the step size it is necessary to learn about the standard deviation of our input and output signals. By getting the step size too big, we are considering long term behavior, but might be missing out on a short term one.
In the near future I will be learning about the meanings of these variables and their contribution to the over all accuracy of our filters.
Results will be posted.
  562   Wed Jun 25 01:30:19 2008 JohnSummaryIOOFrequency noise after the MC
I made some (very) rough estimates of the contribution made to the noise after the mode cleaner by three sources.

Seismic noise - how much of the signal is due to the mode cleaner compenstaing for seismic disturbance of CARM.
Actuator noise - coil drivers and DAC noise.
MC_F - estimate of MC_F suppressed by the loop gain.
Attachment 1: MC2noise080623.png
MC2noise080623.png
  561   Wed Jun 25 00:35:40 2008 KojiSummaryGeneralOptical Layout on the AP table
I have visited the AP table in order to investigate where we are going to put the optical setup for the abs. length meas.
I have attached the PNG and PDF files to share the optical layout. It is not complete. Any comments or corrections are welcome.
Attachment 1: optical_layout_ap_table.png
optical_layout_ap_table.png
Attachment 2: optical_layout_ap_table.pdf
optical_layout_ap_table.pdf
  560   Tue Jun 24 22:43:23 2008 ranaSummarySEIStack TF
Attachment 1: Screenshot.png
Screenshot.png
  559   Tue Jun 24 22:31:10 2008 MashaUpdateAuxiliary lockingmy first step in fiber stabilization
There is a new 1W INNOLIGHT NPRO laser at the Symmetric Port.

Set up an interferometer to measure difference in phase noise introduced by a fiber. Works as expected without the fiber! (balanced intensity, out of phase noise in the two outputs).
  558   Tue Jun 24 17:12:10 2008 josephb, EricConfigurationCamerasGC750 setup, 1X4 Hub connected, ETMX images
The GC750 camera has been setup to look at ETMX. In addition, the new 1X4 rack mounted switch (131.215.113.200) has been connected via new cat6 cable to the control room hub (131.215.113.1?), thanks to Eric. The camera is now plugged into 1X4 rack switch and now has a gigabit connection to the control room computers as well as Mafalda (131.215.113.23).

By using ssh -X mafalda or ssh -X 131.215.113.23, then typing:

target
cd Prosilica/bin-pc/x86/
./Sampleviewer

A viewer will be brought up. By clicking on the 3rd icon from the left (looks like an eye) will bring up a live view.

Closing the view, and then cd ../../40mCode, and then running ./Snap --help will tell you how to use a simple code for taking .tiff images as well as setting things such as exposure length and size of image (in pixels) to send.

When the interferometer was set to an X-arm only configuration, we took two series of 200 images each, with two different exposure lengths.

Attached are three pdf images. The first is just a black and white single image, the second is an average of 100 images, and the third is the standard deviation of the 100 images.
Attachment 1: GC750_ETMX_E30000_single.pdf
GC750_ETMX_E30000_single.pdf
Attachment 2: GC750_ETMX_E30000_avg.pdf
GC750_ETMX_E30000_avg.pdf
Attachment 3: GC750_ETMX_E30000_std.pdf
GC750_ETMX_E30000_std.pdf
  557   Tue Jun 24 15:15:09 2008 JohnSummaryLSCLocking efforts
Rob, Rana, John

In the past week or so we've been working on reducing the CARM offset using a DC signal (SPOB DC).
We were able to get up to arm powers of around 30 (where a single arm cavity lock is a power of 1)
before instability set in and we would lose lock for, as yet, unknown reasons.

In recent nights locking efforts have taken a few backward steps.

Since last Thursday engaging the AO path has proved troublesome, i.e. engaging it would instantly
cause loss of lock. This seems to be related to problems with the mode cleaner servo. For the past
few nights it has been behaving strangely and could not be operated with the usual super boost stages.
Last night the situation was improved. MC boost stages could be used and the AO path engaged. The
cause of this problem and its spontaneous resolution are not understood.

Last night we were unable to switch CARM to SPOB DC. I've attached a spectrum of the MC2 length signal.
This path is being used for CARM and so gives an indication of the frequency noise after the mode
cleaner. At the moment the plot is calibrated in units of Rana's gut feeling. We already tested to see if
any of the excess noise was introduced by the WFS. No evidence was found. We'll try to make a useful
calibration soon and see if our problems are related to excess frequency noise.

Another realisation from last night was the effect of arm detuning on the analogue CARM path. When CARM is detuned
the coupled cavity pole removes an extra 90 degrees of phase. The digital path has the `moving zero' to compensate
for this. The analogue path has no such compensation and can therefore become unstable at moderate detunings.
We propose trying to reduce the CARM offset further before engaging the analogue path. This will give higher
gain and move the UGF to a region of increased phase margin.
Attachment 1: mcl080623.png
mcl080623.png
  556   Tue Jun 24 10:24:43 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Cavity Swing Measurement (2)
At the entry 555, Alberto reported the results of the cavity length measurement using cavity sweeping.
As expected, each result inevitably has an ambiguity depending on which resonance do we take as an upper sideband.

In order to exclude this ambiguity Steve and Koji performed a primitive non-optical measurement using a tape and photos:
This morning Steve and Koji did tape measurements to know the lengths between the ITM/ETM chambers.
Yesterday, Koji took photos of the optical tables in vacuum to know the actual positions of the suspensions.

The results are shown in the figures attached. From those non-optical measurements the lengths of the X/Y arm are known to be 38.48+-0.03 / 38.67+-/0.03 [m].

Then, we could exclude the shorter lengths of the values in the entry 555. i.e. The Y arm is longer than the X arm about 0.2 m.

These approximate lengths will be used in the further precise measurements which use precise scans of the FSR frequencies.
Attachment 1: armlength.png
armlength.png
Attachment 2: armlength2.png
armlength2.png
  555   Mon Jun 23 21:51:19 2008 AlbertoUpdateGeneralArm Cavity Length Measurement
We measured the arm cavity lengths sweeping the ETM mirror position and looking at the reflected demodulated output. We excited the mirror by a sine wave of 0.2 Hz and amplitude of 30000 counts. From the time series of the occurrences of the resonances of the sidebands and of the carrier we evaluated the free spectral range of the cavities and thus the lengths. The details of the procedure are explained in the attached document. As discussed in it, for each cavity we obtain two possible values of the length depending on which of the sideband resonances is that corresponding to the upper sideband and which corresponds to the lower one instead. The numbers are:
Lx=(38.30 +/- 0.08)m / (38.45 +/- 0.08)m
Ly=(38.16 +/- 0.08)m / (38.70 +/- 0.08)m

Since the difference between the two possibilities is quite large, we should be able to decide which one is correct by somehow measuring directly the cavity length. We want to try it tomorrow by a tape meter.


Alberto and Koji
Attachment 1: 40mLengthMeasure.pdf
40mLengthMeasure.pdf 40mLengthMeasure.pdf
  554   Mon Jun 23 19:48:28 2008 rana,albertoSummaryIOOStochMon trends (80 days)
Here's a StochMon plot showing the RFAM after the MC. Remember that in these units, 2V means no RFAM
and 0 V means lots of RFAM. Alberto says "the calibration is in Tiramisu". So there you go.
Attachment 1: e.png
e.png
  553   Mon Jun 23 19:33:01 2008 rana, jenneUpdateIOOMC_F Noise check
We looked at the MC_F spectrum because Rob and Yoichi said that it had gone 'all crazy'. It
seemed fine as we looked at it (even with only one boost stage on) so we looked for things
that might be marginal and make it go nuts.

At the error point (Q mon on the demod board and TEST IN1) of the MC Servo board we saw the
old 3.7 MHz signal (comes from the 33 MHz RFAM getting demodulated by the 29.5 MHz MC LO)
and thought that this might cause some worries. So we installed Jenne's passive elliptic
low pass which has a 3.7 MHz zero.

This wiped out the 3.7 MHz noise but we were not able to re-create the extra frequency noise
so its unlikely to have fixed the main problem. However, we leave it in because its good. If
there is a need to revert it, we have left hanging on the side of the rack the old cable which
was a SMA->TNC making a direct, unfiltered connection between the MC Demod board and the MC
servo board.

More before and after results from Jenne tomorrow, but for now here is a calibrated MC_F spectrum
using the new MC_F-Reference.xml template file.

We also noticed that we could make some small effects on the MCF spec by adjusting the PMC gain so
there's probably more hay to be made there using a lead brick and a gain slider. More in Jenne's
entry.
Attachment 1: mcf.png
mcf.png
  552   Mon Jun 23 15:22:04 2008 ranaBureaucracySAFETYLaser Safety Walkthrough today
Attachment 1: Walkthrough08.jpg
Walkthrough08.jpg
  551   Sun Jun 22 21:38:49 2008 robHowToGeneralIFO CONFIGURE

Now that we're getting back into locking, it's nice to have a stable alignment of the interferometer.
Thus, after you're done with your experiment using subsets of the interferometer (such as a single arm),

please use the IFO_CONFIGURE screen, and click "Restore last Auto-Alignment" in the yellow "Full IFO" section.

If you don't know what this means/how to do this, you shouldn't be using the interferometer on your own.
  550   Fri Jun 20 17:42:48 2008 steveUpdatePSLSS trap scattering compared to black glass trap
Circular SS304 trap was compared to wedged black glass trap.

The measurement set up of entry 529 was changed to define polarization.
CrystaLaser was remounted in horizontal position and half wave plate was placed after it.
These measurements were done in horizontal polarization.

atm1: the cSSt and wBGt was moved horizontally, ~90 degrees of the incident beam
atm2: traps were rotated around the incident beam, ~20 degrees each direction
atm3: set up
atm4: top view of traps
atm5: side view of trap
Attachment 1: cSSt_1.png
cSSt_1.png
Attachment 2: cSSt_2.png
cSSt_2.png
Attachment 3: scatset_61908.png
scatset_61908.png
Attachment 4: cSSt&wBGt.png
cSSt&wBGt.png
Attachment 5: cSStbeam.png
cSStbeam.png
  549   Fri Jun 20 08:30:27 2008 stivUpdatePhotos40m summer line up 2008
atm1: John, Alberto, Yoichi, Koji, Masha, and Sharon

atm2: surf students Max of CIT, Sharon of MIT, Masha of Harvard, Eric of CIT not shown
Attachment 1: P1020559.png
P1020559.png
Attachment 2: P1020560.png
P1020560.png
  548   Fri Jun 20 02:20:33 2008 YoichiUpdate PMC transmittance degradation
The PMC transmitted light power has been degrading constantly for last two weeks (see the attachment 1).
I went down to 2.55V.
The output of the MOPA is constant during this period. More strangely, the reflected power from the PMC is also constant.
One possible explanation is the contamination of the PMC mirrors. But I don't know why it started two weeks ago.

I tweaked the alignment of PMC and was able to recover the transmitted power to above 2.7V (attachment 2).
I will keep eye on this issue.
Attachment 1: pmc_trans_long_trend.pdf
pmc_trans_long_trend.pdf
Attachment 2: pmc_trans_improvement.pdf
pmc_trans_improvement.pdf
  547   Fri Jun 20 01:38:55 2008 ranaUpdatePEM20 day Weather
Yoichi showed me that its possible to make PNG images from PS using GS:
gs -sDEVICE=png16m -sOutputFile=foo.png bar.ps
Attachment 1: test.png
test.png
  546   Thu Jun 19 20:22:03 2008 ranaUpdateGeneralFE Computer Status
I called Rolf (@LLO) who called Alex (@MIT) who suggested that we power cycle every crate
with an RFM connection as we did before (twice in the past year).

Rob and I followed Yoichi around the lab as he turned off and on everything. There
was no special order; he started at the Y-end and worked his way into the corner and
finishing at the X-End. Along the way we also reset the 2 RFM switches around fb0.

This cured the EPICS problem; the FEs could now boot and received the EPICS data.

However, there are still some residual channel hopping-ish issues which Rob and Yoichi are
now working on.
  545   Thu Jun 19 15:52:06 2008 AlbertoConfigurationComputersMeasure of the current absorbed by the new Megatron Computer
Together with Rich Abbot, sam Abbot and I measured the current absorbed by the new Megatron computer that we installed yesterday in the 1Y3 rack. The computer alone absorbs 8.1A at the startup and then goes down to 5.9A at regime. The rest of the rack took 5.2A without the computer so the all rack needs 13.3 at the startup and the 11.1A.

We also measured the current for the 1Y6 rack where an other similar Sun machine has been installed as temporary frame builder and we get 6.5A.


Alberto, Rich and Sam Abbot
  544   Wed Jun 18 18:50:09 2008 ranaUpdateComputersIt can only be attributable to human error. (HAL - 2001)
There has been another one of "those" events and all of the front end machines are down.

We poked around and Rob determined that the FEs can't get the EPICS data from EPICS. The
dcuepics machine is hooked up and running and all of the epics binaries are running. We also
tried resetting its RFM switch as well as power cycling the box using the "poweroff" command.


Not a sausage.

Rob points out that although the Signal Detect lights are on on the cards, the 'Own Data' light
is not on on the dcuepics' card although it is on for some of the cards on the other boxes.


We have placed messages with the Russian. If anyone sees him, don't let him go without fixing things.
Also, make sure to follow him around with notepad and possibly a camera to record what it is that
he does. If he's muttering, maybe try to use a sensitive hidden sound recorder.
  542   Wed Jun 18 18:32:09 2008 Max JonesUpdateComputer Scripts / ProgramsNB Update
I am reconfiguring the the noisebudget code currently in use at the sights. To that end, I have done the following things (in addition to the elog I posted earlier)

In get_dtt_dataset.m - I added C1 specific cases for DARM_CTRL, SEIS, and ITMTRX changing the specific channels to make those in use at caltech

In LocalizeSite.m - I changed the NDS_PATH to match caltechs. I left NDS_HOST untouched.

Since I am trying to get SEIS and DARM to work initially I added C1 specific cases to both of these.

Better documentation may be found in /users/mjones/DailyProgressReport/06_18_08. Suggestions are appreciated. Max.
  541   Wed Jun 18 18:26:19 2008 YoichiUpdatePSLFinding the optimal operation temperature for the NPRO by the slow act scan
Being suspicious of the temperature stabilization of the NPRO crystal, I ran the slow scan script written by Rana to find the suitable operation temperature.
The procedure is the same as the one explained in the entry below:
http://www.ldas-sw.ligo.caltech.edu/ilog/pub/ilog.cgi?group=40m&task=view&date_to_view=09/04/2006&anchor_to_scroll_to=2006:09:04:22:23:56-rana
The attached plots show the results. By looking at C1:PSL-126MOPA_126MON, I set the slow slider voltage to 0.
This time, it looks like the temperature control of the NPRO crystal is working fine.
Obviously, PMC picks up many higher order modes. I will try to mode match/align the PMC later.
Attachment 1: FSS-slow-scan.pdf
FSS-slow-scan.pdf
  540   Wed Jun 18 18:20:10 2008 YoichiUpdatePSLInvestigation on the NPRO temperature stabilization glitches
As Rob pointed out in http://dziban.ligo.caltech.edu:40/40m/537 the MOPA NPRO has been showing some glitchiness in the LTEC loop.
Following Rana's suggestion, Steve and I opened the MOPA and directed a heat gun for a minute to the NPRO hoping that we can see something in the LTEC loop.
The first attachment shows the behavior of LTMP and LTECH along with DTMP and DTECH at the time of the heat gun attack.
T=0 is the time when Steve directed the heat gun to the NPRO. There is no response neither in LTMP nor LTECH.
DTMP and DTECH look like responding.
Around the center, there is a dip in LTMP. This might be caused by removing the heat gun. But we are not sure. This kind of small glitches can be found in LTMP everywhere (see the attachment 2).
It looks like the LTMP sensor is not working, or the LTECH loop is actually working but the LTECH reading is broken.
However, the scan of the slow actuator (temperature) shows the LTECH loop is actually working. So it is a bit confusing.
More investigation is necessary.
See the next entry by me.
Attachment 1: LTEC-loop-HeatGun-Response.pdf
LTEC-loop-HeatGun-Response.pdf
Attachment 2: LTMP-glitches.pdf
LTMP-glitches.pdf
  539   Wed Jun 18 16:37:54 2008 steve,ranaUpdateSAFETYCO2 test in the east arm
The CO2 laser and table are in the east arm for characterization of the mechanics. We
will not be operating it until we have an SOP (which is being written). No worries.
Attachment 1: co2.png
co2.png
  538   Wed Jun 18 16:07:57 2008 robSummaryComputersRFM network down

The RFM network tripped off around noon today. It's still down. The problem appears to be with the EPICS interface (c1dcuepics). Trying to restart one of the end stations yields the error: No response from EPICS.

Possible causes include (but not limited to): busted RFM card on c1dcuepics, busted PMC bus on c1dcuepics, busted fiber from c1dcuepics to the RFM switch. We need Alex.
ELOG V3.1.3-